Given that P(A) =0.55 P(B) =0.27 and P(A and B) = 0.13 if events A and B are mutually exclusive find P(A or B)
1
Expert's answer
2019-09-09T12:29:48-0400

For two mutually exclusive events


"P(A\ or\ B)=P(A)+P(B)
\\P(A\ or\ B)=0.55+0.27=0.82"